OpenStudy (anonymous):

dy/dx = (cos x) e^(y+sinx) and y = 0 when x = 0. find the original equation

OpenStudy (cgreenwade2000):

Original equation as in the integral of that function?

OpenStudy (anonymous):

as an answer I get e^sinx + ln1 -1 =y which is not the right answer, according to the book

OpenStudy (anonymous):

yes sir, we are to integrate via separation of variables

OpenStudy (cgreenwade2000):

This website can help faster than I can. http://tutorial.math.lamar.edu/Classes/DE/SeparationofVariables.aspx Where I went with all of my questions.

OpenStudy (anonymous):

maybe you can check my work \[(\cos x)(e ^{y+\sin x})\] \[ \frac{ dy }{ e^y } = (\cos x) e^y dx\] u = sinx and du = cos x \[\ln e^y + c = e^{sinx} +c\] find c by plugging in 0 and I get (ln 1) -1 = c put it all together and I get e^[sinx] + ln 1 - 1 = y

OpenStudy (anonymous):

First you must separate the variables. \[\frac{ dy }{ dx } = cosx e ^{y+sinx}\] \[\frac{ dy }{ dx }=cosx e ^{y}e ^{sinx}\] \[\int\limits_{?}^{?}\frac{ dy }{ e ^{y} }=\int\limits_{?}^{?}cosxe ^{sinx}\] for the left part of the equation, let u=sinx then du=cosxdx I think you can do the rest. Cheers

OpenStudy (anonymous):

actually, its after that where i begin to get stuck, can you go a little further? no need to use the equations button, just a verbal explanation will suffice

OpenStudy (cgreenwade2000):

For the right side, let u = sinx, then du = cosxdx

OpenStudy (anonymous):

so substitute u and du, this will result to the equation \[\int\limits_{?}^{?}=e ^{u}du\] cos x will be cancelled, after the integration, you will com up to \[e ^{u}\] wich is equal to \[e ^{sinx}\]

OpenStudy (anonymous):

ln(e^y)+c=e^sinx+c what happens after this?

OpenStudy (anonymous):

integral of 1/ e^y is ln e^y isn't it?

OpenStudy (cgreenwade2000):

On the left side, substitute u =-y and then -du=dy

OpenStudy (anonymous):

ya, I was wrong. you got it, just substitute the constraints to find C

OpenStudy (cgreenwade2000):

1/e^y = e^-y

OpenStudy (anonymous):

holy cow, thats what i needed to see haha

OpenStudy (anonymous):

wait, I think I was correct

OpenStudy (cgreenwade2000):

Then substitute. Giving you the (negative) integral e^u du

OpenStudy (anonymous):

yeah jeffrey i think you were right, I just did not get that you turned 1/e^y to e^-y as wade just showed me

OpenStudy (anonymous):

ya, so therefore the final equation would be \[-\frac{ 1 }{ e ^{y} }=e ^{sinx} +C\]

OpenStudy (cgreenwade2000):

Yes....

OpenStudy (anonymous):

There you go, substitute the constraints to find C.

OpenStudy (anonymous):

alright sweet, wish i could give you both best answer, seriously, thanks to both of you

OpenStudy (anonymous):

No prob dude! :)